What is the percent yield of AlBr3 if 120.7 grams of AlBr3 is produced in when 12.7 g of aluminum reacts? Round to two decimal places and do not include a % sign in your answer.


2Al + 3Br2 → 2AlBr3
1
Expert's answer
2019-03-12T08:55:00-0400

2Al + 3Br2"\to" 2AlBr3

theoretical yield = (m(Al)*2*M(AlBr3))/2*M(Al)="(12.7g*2*266.69g\/mol)\/(2\u221726.98g\/mol)=125.536g"

%yield=actual yield/theoretical yield*100%="120.7g\/125.536g*100 =96.15"